What will be the quantitative concentration of the newly formed solution if we add 800 cm3 of water to 400 cm3 of hydrochloric acid solution, mass concentration of 12.25 g / dm3?


1
Expert's answer
2021-05-11T05:35:09-0400

Solution:

hydrochloric acid - HCl

Convert units:

(800 cm3) × (0.001 dm3 / 1 cm3) = 0.8 dm3 H2O

(400 cm3) × (0.001 dm3 / 1 cm3) = 0.4 dm3 HCl


(0.4 dm3 HCl) × (12.25 g HCl / 1 dm3 HCl) = 4.9 g HCl

New solution volume = 0.8 dm3 + 0.4 dm3 = 1.2 dm3

Hence,

(4.9 g HCl / 1.2 dm3) = 4.08 g/dm3 - the new mass concentration of HCl solution

OR:

(4.08 g/dm3) × (1 mol / 36.458 g) = 0.112 mol/dm3 = 0.112 M - the new molarity of HCl solution


Answer: 4.08 g/dm3 or 0.112 M is the quantitative concentration of the newly formed solution.